The result of nationwide aptitude test in mathematics are normally distributed with m=80 and s= 15 what is the percentile rank of a score of 84


1
Expert's answer
2022-02-15T14:00:09-0500

"\mu=80\\\sigma=15"

We first determine, "p(x\le 84)=p({x-\mu\over \sigma}\le{84-\mu\over \sigma})=p(Z\le {84-80\over 15})=p(Z\le0.27)=0.6064"

The percentile rank is, "0.6064\times100\%=60.64\%"

Therefore, with score, "x=84", the percentile rank is 60.64 %